HETAV PATEL

Grade 11,

∫ dx/x 2 -1 In the above equation before - one is x square.


∫ dx/x2-1




In  the above equation before - one is x square.



 

Grade:12th Pass

3 Answers

Chetan Mandayam Nayakar
12 years ago

1/(x^2 -1)=1/(x+1)(x-1)=(1/2)((1/(x+1))-(1/(x-1)))

Answer=(1/2)ln|(x+1)/(x-1)|

yeshwendra shukla
12 years ago

1/(x^2 -1)=1/(x+1)(x-1)=(1/2)((1/(x-1))-(1/(x+1)))

Answer=(1/2)ln|(x-1)/(x+1)|

 

this one is the right answer dude.....well all the best...:)

Subhojit Paul
12 years ago

integration of 1/(1-x2) is tan-1 x.

Thus here in this problem, if we take a minus sign common,then we will get the equation in this form. thus the answer is

-tan-1x